# Score-based Generative Priors Guided Model-driven Network for MRI Reconstruction

### v1

@@ V-C Ablation Studies
TABLE II: Results of ablation studies on coronal PD sequence with four- and six-fold acceleration rates.
@@
SG 26.8198$\pm$4.7495 25.0439$\pm$7.2973 0.4172$\pm$0.1712 0.5212$\pm$0.1568
MN 38.8362$\pm$2.6402 35.9488$\pm$2.6333 0.8623$\pm$0.0880 0.7861$\pm$0.1088
SR-MN 35.4585$\pm$2.9562 33.2093$\pm$2.6398 0.8351$\pm$0.0912 0.7412$\pm$0.1284
SG-MN-DM-US 40.7095$\pm$3.0165 38.0227$\pm$2.9647 0.9619$\pm$0.0615 0.9440$\pm$0.0700
SG-MN-DM-DG 40.8443$\pm$3.0384 38.2183$\pm$2.9986 0.9629$\pm$0.0624 0.9448$\pm$0.0721
SG-MN-US-DG 41.0424$\pm$3.4144 38.1616$\pm$2.9985 0.9633$\pm$0.0676 0.9443$\pm$0.0714
SG-MN-DM-US-DG 41.2213$\pm$3.3385 38.4681$\pm$2.9889 0.9638$\pm$0.0642 0.9459$\pm$0.0678
To highlight the improvements brought by different modules, we designed several workflows for a clear evaluation. Specifically, the proposed modules for evaluation were the SMLD-based guidance (SG), model-driven network structure (MN), denoising module (DM), updating steps (US) for SGs, and densely connected guidance (DG) in the unrolled network. The ablations was designed as follows.
@@
    MN: a standard unrolled network with the undersampled measurements as the inputs.
@@
  5. 5.
    SG-MN-DM-DG: the SMLD results are used as guidance in the unrolled network but without and updating steps.
@@
    SG-MN-US-DG: the SMLD results are used to guide the network without denoising in the DM.
@@
The quantitative results are shown in table II. Comparing SG with the proposed SGM-Net, we can observe a significant improvement brought about by the proposed modules, whereas the average results from the SMLD sampling are worse than those of the zero-filled method. Similarly, through the guidance of the SMLD results, the results of the unrolled MN method were significantly improved with guidance from the SMLD results. Using either of these two methods alone can’t achieve satisfactory results. Moreover, we simply combine SR and MN in the SR-MN workflow that the SMLD results are directly fed to the network as input. However, the performance are even worse than that of MN. This phenomenon demonstrates the effectiveness of the proposed guidance-based workflow, which outperforms a simple combination of the two methods.
@@
A comparison between SG-MN-DM-US and SGM-Net demonstrated that the densely connected structure is capable of fulfilling the guiding work and improving the final results. Removing the densely connected structure and using only the AMs to optimize the reconstructed images will lead to a degredation in the final results. By comparing SG-MN-DM-DG with SGM-Net, we observed that the updating steps in the proposed model can ensured that accurate guidance was provided by the PGIs and improved the results. When removing the DM and directly feeding SMLD results to guide the network, the reconstruction performances of SG-MN-US-DG are similarly worse than that of SGM-Net. The best results are achieved with the integration of all the proposed modules, which demonstrates their effectiveness and necessity. Furthermore, we illustrate in subsection V.E that the integration of the proposed module can enhance the robustness of the network with greatly reduced training data.
Figure 5: The reconstruction results of the original model at four-fold acceleration. The first rows of each subfigure are the reconstructed images, the second rows are zoomed details in the red square, and the third rows are the error maps representing the objective reality. The corresponding images from left to right: the objective reality, zero-filled measurements, PGI, DGI, and the final output of $x_{z}^{k}$ and $x_{T}^{k}$ branch.
@@
Figure 6: The reconstruction results with 20% sampling steps at four-fold acceleration. The first rows of each subfigure are the reconstructed images, the second rows are zoomed details in the red square, and the third rows are the error maps representing the objective reality. The corresponding images from left to right: the objective reality, zero-filled measurements, PGI, DGI, and the final output of $x_{z}^{k}$ and $x_{T}^{k}$ branch.
@@
TABLE III: Quantitative results of the selected images at four-fold and six-fold acceleration rates.
@@
zero-filled 31.3089$\pm$3.3939 30.5176$\pm$3.3902 0.8778$\pm$0.0713 0.8626$\pm$0.0799
original $x_{T}$ 26.8198$\pm$4.7495 25.0439$\pm$7.2973 0.4172$\pm$0.1712 0.5212$\pm$0.1568
$x_{T}^{0}$ 39.0871$\pm$2.5214 37.0431$\pm$2.3824 0.9523$\pm$0.0587 0.9364$\pm$0.0635
$x_{T}^{K}$ 41.0913$\pm$3.4621 38.3504$\pm$2.9553 0.9635$\pm$0.0656 0.9458$\pm$0.0678
$x_{z}^{K}$ 41.2213$\pm$3.3385 38.4681$\pm$2.9889 0.9638$\pm$0.0642 0.9459$\pm$0.0678
20% sampling steps $x_{T}$ 19.7481$\pm$7.8121 17.5994$\pm$7.7102 0.4664$\pm$0.2049 0.4312$\pm$0.1723
$x_{T}^{0}$ 38.4951$\pm$2.9590 36.4403$\pm$2.9405 0.9500$\pm$0.0600 0.9317$\pm$0.0656
$x_{T}^{K}$ 40.8341$\pm$3.8591 38.2336$\pm$3.0403 0.9631$\pm$0.0654 0.9421$\pm$0.0691
$x_{z}^{K}$ 40.9907$\pm$3.7419 38.3414$\pm$2.8912 0.9631$\pm$0.0657 0.9446$\pm$0.0682
pretrained with another dataset $x_{T}$ 25.5783$\pm$4.7741 24.2301$\pm$7.8034 0.4904$\pm$0.1506 0.4692$\pm$0.1549
$x_{T}^{0}$ 39.0112$\pm$2.7413 36.8603$\pm$2.7916 0.9510$\pm$0.0578 0.9351$\pm$0.6225
$x_{T}^{K}$ 40.9581$\pm$3.3235 38.2714$\pm$2.8405 0.9634$\pm$0.0663 0.9401$\pm$0.0702
$x_{z}^{K}$ 41.0920$\pm$3.2329 38.4204$\pm$2.9084 0.9636$\pm$0.6828 0.9432$\pm$0.0688
V-D Evaluation of the guided reconstruction process
To evaluate the improvements brought by the guided process, we plotted the following images: the image of the ground truth, zero-filled measurements, PGI ($x_{T}$), DGI ($x_{T}^{0}$), and the final output of $x_{z}^{k}$ and $x_{T}^{k}$ branches ($x_{T}^{K}$ and $x_{z}^{K}$) of an MR image from the test dataset. To mitigate potential influences from the sampling process, we designed the following three sampling methods:
  1. 1.
    original: the SMLD results are sampled using the pretrained score network, as described previously.
  2. 2.
    20% sampling steps: the SMLD results are sampled using the pretrained score network, but using only 20% of the original sampling steps.
  3. 3.
    pretrained with another dataset: we continued to train the score network on another brain dataset proposed in [69], and the images are sampled with the same hyperparameters in the original method.
@@ V-D Evaluation of the guided reconstruction process
The reconstructed images from the original sampling method at four-fold acceleration are shown in Fig. 5. The first row shows the corresponding images, and the second row shows the magnified details. The zero-filled image (second column) is blurred, and the PGI (third column) loses texture details. However, the artifacts in the DGI (fourth column) were effectively eliminated from the zero-filled image and PGI by DMs. Then, the GICs recovered the missing details in the DGIs and $x_{T}^{K}$ (fifth column) and $x_{z}^{K}$ (sixth column) presented satisfactory results when compared with the ground truth images (first column). The error maps in the third row exhibit that the massive errors in the zero-filled images and PGIs are obviously alleviated, the the image quality are further improved in $x_{T}^{K}$ and $x_{z}^{K}$. The reconstructed images with 20% sampling steps at four-fold acceleration are depicted in Fig. 6, which are also consistent with the observations in Fig. 5.
@@
The quantitative results for the three sampling methods are presented in Table III. $x_{z}^{K}$, which is chosen as the final result, slightly outperforms $x_{T}^{K}$, and the other results obtained at four- or six-fold acceleration are consistent with the results observed from the images. Significantly reducing the sampling steps leads to further deterioration in the quality of SMLD results. However, the impact on the final reconstruction results is minimal, demonstrating the robustness of our proposed method to noise input, and the DMs are effective to extract valuable information from the noise input. We continue training the score network on another brain dataset to make perturbations on the pretrained model, which produces reconstructions of equally high quality through our workflow, demonstrating that our method imposes minimal requirements on the sampling process. In summary, this experiment vividly illustrates how our proposed workflow leverages SMLD results to extract crucial information, aiding in the reconstruction process and finally achieving high-quality results.
V-E Evaluation of the network robustness
@@
Figure 7: PSNR (the first column) and SSIM (the second column) results of SGM-Net, MeDL-Net, and E2EVarNet with different amounts of training data at four- (the first row) and six-fold (the second row).
In this subsection we illustrate the network robustness with limited training data. Here, we compared E2E-VarNet and MeDL-Net, which have convincing performances and do not leverage guidance images, as shown in Table I. We also compared with three ablation networks (i.e. SG-MN-DM-US, SG-MN-MN-DG and SG-MN-US-DG). We reduced the number of training images of the different methods to different extents and compared their performances with our model. The results are presented in Fig. 7. We compared the PSNR and SSIM results of the above networks at four- and six-fold, respectively. We observed that the performances of the other networks deteriorated with decreasing training data. Meanwhile, the performance of SGM-Net was more stable even with only 20 training images. Compared with E2E-VarNet and MeDL-Net, our ablation networks also exhibit relatively better robustness. However, the degradation of image quality become more pronounced with the decrease of training data, which further demonstrating the necessity of the proposed modules. To sum up, this experiment demonstrates that proposed guidance process can effectively compensate for information in the missing training data and make the model robust.
VI Discussions
@@
To make the sampling process robust and achieve better performance, most of the cutting-edge methods proposed to optimize the training of the generative models or the sampling steps. For example, the authors in [14] fed high-dimensional tensors to a noise conditional score network during training, and they designed homotopic iterations at the inference phase. The authors in [13] proposed a framework named score-POCS. During inference, the reconstruction process iterates between SDE solvers and data consistency steps. Compared with them, the most significant distinction in our workflow is that we bypass all the tedious hyperparameter tuning and network retraining steps and implement a naive sampling. The reconstruction performances are then not robust especially with distribution-shift test datasets, and the results are observed to be severely corrupted. We novelly treat the SMLD results as a generalized prior, and we design an unrolled network which can extract valuable information from the SMLD samples to guide reconstruction, while the components during sampling can also help to eliminate the artifacts in the network. Experimental results have demonstrated that compared with the naive sampling results, only small amounts of data are required to train the unrolled network and the performances can be greatly improved.
@@
Another bottleneck in diffusion models is that the sampling process is relatively slow. To solve this problem, the authors in [70] proposed DiffuseRecon, and the reverse process can be guided with observed k-space signal. The authors designed a coarse-to-fine algorithm to save sampling time. In [71], the authors proposed deep ensemble denoisers and incorporated them with score-based generative models to reconstructed images. Consequently, the necessary iterations are reduced. In our workflow, we preserved the naive sampling process. However, the unrolled network have good tolerance to artifact-affected inputs, hence the sampling steps can also be great reduced.
@@
In the proposed reconstruction workflow, the low-quality SMLD samples have been demonstrated to be effective to guide the network training and improve the final results. Meanwhile, the proposed workflow is also compatible to optimized SMLD steps with higher quality references, which show great potential to be exploited in future works. Besides, a limitation of the proposed method is that although the workflow exhibits considerable tolerance for the SMLD process, the unrolled network still requires some data for training to achieve superior performance. Training strategies such as self-supervising should be considered to further optimize the workflow in future works.
VII Conclusions

### v2

@@ 5.3 Ablation Studies
Table 2: Results of ablation studies on coronal PD sequence with four- and six-fold acceleration rates.
@@
SG 31.3478$\pm$2.6593 29.7293$\pm$5.1889 0.4730$\pm$0.1556 0.6067$\pm$0.1215
MN 38.2080$\pm$2.4697 35.8126$\pm$2.4974 0.8346$\pm$0.1386 0.7849$\pm$0.1279
SG-MN 35.4585$\pm$2.9562 33.2093$\pm$2.6398 0.8351$\pm$0.0912 0.7412$\pm$0.1284
SG-MN-DM 39.3722$\pm$3.6295 36.9837$\pm$4.0466 0.9533$\pm$0.0801 0.9368$\pm$0.0812
SG-MN-DM-US 40.5941$\pm$3.0464 37.8509$\pm$2.8824 0.9615$\pm$0.0619 0.9425$\pm$0.0684
SG-MN-DM-DG 41.0969$\pm$3.7561 38.3457$\pm$2.9926 0.9625$\pm$0.0803 0.9440$\pm$0.0730
SG-MN-US-DG 41.1213$\pm$3.3004 38.3419$\pm$3.0769 0.9630$\pm$0.0715 0.9444$\pm$0.0772
SG-MN-DM-US-DG 41.1981$\pm$3.3802 38.4491$\pm$3.2166 0.9634$\pm$0.0682 0.9452$\pm$0.0790
Table 3: Reconstruction results on coronal PD sequence with four- and six-fold acceleration rates using different inputs in DMs.
structure Method PSNR SSIM
w/o unrolled network $x_{T}$ 38.5337$\pm$3.8373 36.2220$\pm$4.5657 0.9487$\pm$0.0812 0.9322$\pm$0.0825
$x_{T},x_{d}$ 38.9168$\pm$3.9265 36.4447$\pm$4.5595 0.9496$\pm$0.0808 0.9327$\pm$0.0818
$x_{T},x_{p}$ 38.9123$\pm$3.4768 36.6218$\pm$4.1896 0.9507$\pm$0.0799 0.9345$\pm$0.0814
$x_{T},x_{d},x_{p}$ 39.3722$\pm$3.6295 36.9837$\pm$4.0466 0.9533$\pm$0.0801 0.9368$\pm$0.0812
w unrolled network $x_{T}$ 41.1164$\pm$3.4041 38.2711$\pm$3.6271 0.9629$\pm$0.0708 0.9441$\pm$0.0825
$x_{T},x_{d}$ 41.0953$\pm$3.3920 38.2932$\pm$3.3948 0.9628$\pm$0.0669 0.9438$\pm$0.0814
$x_{T},x_{p}$ 41.1620$\pm$3.1978 38.4075$\pm$3.3948 0.9633$\pm$0.0639 0.9450$\pm$0.0812
$x_{T},x_{d},x_{p}$ 41.1981$\pm$3.3802 38.4491$\pm$3.2166 0.9634$\pm$0.0682 0.9452$\pm$0.0790
To highlight the improvements brought by different modules, we designed several ablation workflows for a clear evaluation. Specifically, the proposed modules for evaluation were the SMLD-based guidance (SG), model-driven network structure (MN), denoising module (DM), updating steps (US) for SGs, and densely connected guidance (DG) in the unrolled network. The ablations was designed as follows.
@@
    MN: a standard unrolled network without guidances. The undersampled measurements are adopted as the inputs.
@@
    SG-MN-DM-DG: the SMLD results are used as guidance in the unrolled network but without the updating steps.
@@
    SG-MN-US-DG: the SMLD results are used to guide the network without denoising by the DM.
  8. 8.
@@
The quantitative results are shown in table 2. Comparing SG with the proposed SGM-Net, we observed a significant improvement brought by the proposed modules, whereas the average results from the SMLD sampling were even worse than those of the zero-filled method. Similarly, through the guidance of the SMLD results, the results of the unrolled MN method were significantly improved. Using either of these two methods alone could not achieve satisfactory results. Moreover, we simply combined SG and MN in the SG-MN workflow that the SMLD results were directly fed to the network as input. However, the performance was worse than that of MN. This phenomenon demonstrated the effectiveness of the proposed guidance-based workflow, which outperformed a simple combination of the two methods.
Comparing with the proposed SGM-Net, the SG-MN-DM workflow removed the guided reconstruction structures, and we observed deterioration on the results. On the other hand, the improvements brought by the DMs over SG and MN method were also obvious.
@@
A comparison between SG-MN-DM-US and SGM-Net demonstrated that the densely connected structure was capable of fulfilling the guiding work and significantly improving the final results. Removing the densely connected structure and using only the AMs to optimize the reconstructed images would lead to a degredation in the final results. By comparing SG-MN-DM-DG with SGM-Net, we observed that the updating steps in the proposed model can further ensured that accurate guidance was provided by the PGIs and improved the results. When removing the DM and directly feeding SMLD results to guide the network, the reconstruction performances of SG-MN-US-DG were similarly worse than that of SGM-Net.
Meanwhile, when removing DM or US structure from SGM-Net, we observed that the network still achieved convincing results. Specifically, when removing US, the denoised output of DM can provide effective guidance for the subsequent reconstructions. When removing DM, the updating steps in the networks improved the image quality of SMLD results during training. This phenomenon further demonstrated the robustness of the proposed network structure. The best results were achieved with the integration of all the proposed modules, which demonstrated their effectiveness and necessity. Furthermore, we illustrate in subsection 5.5 that the integration of the proposed module can enhance the robustness of the network with greatly reduced training data.
Besides, we conducted experiments using different features in DMs, and the results are shown in Table 3. Here, we evaluated the results with and without the guided reconstruction process. When the unrolled network structure was removed, the improvements brought by involving $x_{d}$ or $x_{p}$ in DMs is evident, which clearly exhibited the improvements brought by the SIE and CIE blocks. With the unrolled network structure, the distances among the methods were narrowed, that the unrolled network can compensate for the deterioration in DMs. However, incorporating all the components in DMs exhibited the best results, demonstrating the necessity of the proposed SIE and CIE blocks.
@@
Figure 5: (a) The reconstruction results of the original model at four-fold acceleration. (b) The reconstruction results with 20% sampling steps at four-fold acceleration. The first rows of each subfigure are the reconstructed images, and the second rows are zoomed details in the red squares. The corresponding images from left to right: the ground truth, zero-filled measurements, PGI, DGI, and the final output of $x_{T}^{K}$ and $x_{z}^{K}$ branch.
Figure 6: (a) The reconstruction results on fastMRI dataset at four-fold acceleration. (b) The reconstruction results on fastMRI dataset at six-fold acceleration. The first rows of each subfigure are the reconstructed images and the second rows are zoomed details in the red squares. The corresponding images from left to right: the ground truth, zero-filled measurements, PGI, DGI, and the final output of $x_{T}^{K}$ and $x_{z}^{K}$ branch.
@@
Table 4: Quantitative results of the selected images at four-fold and six-fold acceleration rates.
@@
original, Coronal PD zero-filled 31.3089$\pm$3.3939 30.5176$\pm$3.3902 0.8778$\pm$0.0713 0.8626$\pm$0.0799
$x_{T}$ 31.3478$\pm$2.6593 29.7293$\pm$5.1889 0.4730$\pm$0.1556 0.6067$\pm$0.1215
$x_{T}^{0}$ 38.9772$\pm$3.6659 37.0103$\pm$4.1097 0.9508$\pm$0.0798 0.9370$\pm$0.0812
$x_{T}^{K}$ 41.1457$\pm$3.6320 38.2893$\pm$3.9790 0.9630$\pm$0.0750 0.9450$\pm$0.0828
$x_{z}^{K}$ 41.1981$\pm$3.3802 38.4491$\pm$3.2166 0.9634$\pm$0.0682 0.9452$\pm$0.0790
20% sampling steps, Coronal PD zero-filled 31.3089$\pm$3.3939 30.5176$\pm$3.3902 0.8778$\pm$0.0713 0.8626$\pm$0.0799
$x_{T}$ 22.5488$\pm$7.6208 20.6821$\pm$7.8592 0.4664$\pm$0.2049 0.4601$\pm$0.2810
$x_{T}^{0}$ 37.8958$\pm$4.2813 36.6910$\pm$4.1992 0.9461$\pm$0.0809 0.9326$\pm$0.0837
$x_{T}^{K}$ 40.9129$\pm$4.2478 38.2178$\pm$3.8602 0.9622$\pm$0.0817 0.9403$\pm$0.0826
$x_{z}^{K}$ 40.9847$\pm$3.6107 38.2610$\pm$3.6829 0.9628$\pm$0.0770 0.9412$\pm$0.0798
original, FastMRI zero-filled 33.2828$\pm$3.0594 32.5069$\pm$3.0657 0.8857$\pm$0.0709 0.8680$\pm$0.0791
$x_{T}$ 26.5182$\pm$5.1008 23.9689$\pm$5.0221 0.4081$\pm$0.1377 0.4533$\pm$0.1307
$x_{T}^{0}$ 37.3427$\pm$3.5911 35.0240$\pm$3.3366 0.9236$\pm$0.0682 0.8933$\pm$0.0799
$x_{T}^{K}$ 38.7848$\pm$3.4290 36.8528$\pm$3.0269 0.9338$\pm$0.0591 0.9092$\pm$0.0705
$x_{z}^{K}$ 38.8241$\pm$3.3808 36.8730$\pm$3.0364 0.9344$\pm$0.0588 0.9094$\pm$0.0707
5.4 Evaluating the effectiveness of the proposed workflow
@@
To evaluate the effectiveness of the proposed workflow, we plotted the following images: the ground truth, zero-filled measurements, PGI ($x_{T}$), DGI ($x_{T}^{0}$), and the final output of $x_{z}^{k}$ and $x_{T}^{k}$ branches ($x_{T}^{K}$ and $x_{z}^{K}$) of an MR image from the test dataset. The reconstructed images from the original sampling method at four-fold acceleration are shown in Fig. 5 (a). The first row shows the corresponding images, and the second row shows the magnified details. We can observe that the zero-filled image (second column) is blurred, and the PGI (third column) exhibits hallucination artifacts. However, the artifacts in the DGI (fourth column) were effectively eliminated by DMs from the zero-filled image and PGI. Then, the GICs further enhanced the image quality of the DGIs, and $x_{T}^{K}$ (fifth column) and $x_{z}^{K}$ (sixth column) presented satisfactory results when compared with the ground truth images (first column), demonstrating the proposed coarse-to-fine workflow has powerful ability in recovering fine details. Besides, we sampled PGIs from the pretrained score network, but using only 20% of the original sampling steps. The reconstructed images with 20% sampling steps at four-fold acceleration are depicted in Fig. 5 (b). Adopting PGIs with worse quality, the proposed method still robustly achieve high-quality results, which are also consistent with the observations in Fig. 5 (a). Besides, we plotted reconstructed images from the FastMRI dataset at four-fold and six-fold in Fig. 6 (a) and (b). We observed unexpected artifacts on the PGIs, which lower the image quality. However, through the denoising steps, we obseved that the artifacts have been greatly suppressed in DGIs, and the outputs of the unrolled network further eliminated the influence of noise and artifacts.
@@
The quantitative results for the abovementioned methods are presented in Table 4. $x_{z}^{K}$, which is chosen as the final result, slightly outperforms $x_{T}^{K}$, and the other results obtained at four- or six-fold acceleration are consistent with the results observed from the images. Significantly reducing the sampling steps leads to further deterioration in the quality of SMLD results. However, the impact on the final reconstruction results is minimal, demonstrating the robustness of our proposed method to noise input, and the DMs are effective to extract valuable information from the noise input. In summary, this experiment vividly illustrates how the coarse-to-fine workflow leverages SMLD results to extract crucial information, aiding in the reconstruction process and finally achieving high-quality results.
5.5 Evaluating the network robustness with limited training slices
@@
Figure 7: Four- and six-fold PSNR and SSIM results of different networks with limited training data.
In this subsection we evaluated the network robustness with limited training data. Here, we first compared with E2E-VarNet and MeDL-Net, which have shown convincing performances without utilizing guidance images. We compared them with SGM-Net on the coronal PD sequence, and the PSNR and SSIM results at four-fold and six-fold are depicted in Fig 5.5 (a)-(d). We trained the networks with 20, 50 and 100 slices, which are randomly selected, to evaluate their robustness. We observed that the performances of E2E-VarNet and MeDL-Net deteriorated with decreasing training data. Meanwhile, the performance of SGM-Net was more stable with greatly reduced training slices.
We also compared with three ablation networks (i.e. SG-MN-DM-US, SG-MN-MN-DG and SG-MN-US-DG), and the results are depicted in Fig 5.5 (e)-(h). Compared with E2E-VarNet and MeDL-Net, the ablation networks similarly showed better robustness, with a slower degradation in image quality under limited training data conditions, which also indicates the robustness of the proposed workflow. Furthermore, the degradation of image quality become more pronounced with the decrease of training data, which demonstrated the necessity of the proposed modules. To sum up, this experiment demonstrated that proposed guided reconstruction process can effectively compensate for information in the missing training data and achieve convincing performance robustly.
6 Discussions
@@
To make the sampling process robust and achieve better performance, most of the cutting-edge methods proposed to optimize the training of the generative models or the sampling steps. For example, the authors in [13] fed high-dimensional tensors to a noise conditional score network during training, and they designed homotopic iterations at the inference phase. The authors in [12] proposed a framework named score-POCS. During inference, the reconstruction process iterates between SDE solvers and data consistency steps. In [15], the authors enhanced the reconstruction robustness of diffusion models with test-time hyperparameter tuning based on Stein’s unbiased risk estimator. Compared with them, the most significant distinction in our workflow is that we bypassed all the tedious hyperparameter tuning and network retraining steps and implemented a naive sampling. The reconstruction performances are then not robust especially with distribution-shift test datasets, and the results are observed to be severely corrupted. We treat the SMLD results as a generalized prior, and we designed an unrolled network which can extract valuable information from the SMLD samples to guide reconstruction. Experimental results have demonstrated that compared with the naive sampling results, only small amounts of data are required to train the unrolled network and the performances can be greatly improved.
@@
Another bottleneck in diffusion models is that the sampling process is relatively slow. To solve this problem, the authors in [73], for example, proposed DiffuseRecon, and the reverse process can be guided with observed k-space signal. The authors designed a coarse-to-fine algorithm to save sampling time. In [74], the authors proposed deep ensemble denoisers and incorporated them with score-based generative models to reconstructed images. Consequently, the necessary iterations are reduced. In our workflow, we preserved the naive sampling process. However, the unrolled network have good tolerance to artifact-affected inputs, hence the sampling steps can also be great reduced.
@@
In the proposed reconstruction workflow, the coarse SMLD samples have been demonstrated to be effective to guide the network training and improve the final results. Meanwhile, the proposed workflow is also compatible to optimized SMLD steps with higher quality references, which show great potential to be exploited in future works. Besides, a limitation of the proposed method is that although the workflow exhibits considerable tolerance for the SMLD process, the unrolled network still requires some data for training to achieve superior performance. Training strategies such as self-supervising should be considered to further optimize the workflow in future works.
7 Conclusions